# Event Topology Classifiers at the Large Hadron Collider

### v1

@@ II.8 Charged-particle flattenicity ($\rho_{\rm{ch}}$)
Isotropic events are expected to have a uniform distribution of $p_{\rm T}$ throughout the ($\eta-\phi$) region. To quantify the uniformity of $p_{\rm T}$ distribution event-by-event, the ($\eta-\phi$) region is divided into a $(10\times 10)$ grid. $p_{\rm T}$ in each cell ($p_{\rm T}^{\rm cell}$) can be measured, and corresponding mean ($\langle p_{\rm T}^{\rm cell}\rangle$) and standard deviation ($\sigma_{p_{\rm T}^{\rm cell}}$) can be estimated. Keeping the ALICE3 tracking capabilities in mind, the charged particles in $|\eta|<4.0$ and $p_{\rm T}>0.15$ GeV/c are considered for the estimation of $\langle p_{\rm T}^{\rm cell}\rangle$ and $\sigma_{p_{\rm T}^{\rm cell}}$. Thus, for an event, flattenicity ($\rho$) is defined as the relative standard deviation as follows [105],
@@
Figure 20: Correlation of flattenicity with number of MPI
Events with both jet signals and underlying event activity are expected to have a larger spread, and correspondingly, $\sigma_{p_{\rm T}^{\rm cell}}$ will be higher. In contrast, events having only soft production of particles will have a narrower distribution of $p_{\rm T}^{\rm cell}$, and therefore corresponding $\sigma_{p_{\rm T}^{\rm cell}}$ will be smaller. It is important to note that the presence of jet signals in the jetty events can also enhance $\langle p_{\rm T}^{\rm cell}\rangle$ in the denominator of Eq. (20); however, the increase in $\langle p_{\rm T}^{\rm cell}\rangle$ will be overshadowed by the corresponding rise in $\sigma_{p_{\rm T}^{\rm cell}}$ in the numerator. Thus, the value of $\rho$ is expected to be smaller for isotropic events, while a large value of $\sigma_{p_{\rm T}^{\rm cell}}$ corresponds to a jetty event.
However, the current PID and tracking capabilities at ALICE (LHC) and the STAR experiment (RHIC) are limited to $|\eta|<1.0$. As a result, using the above definition, one can not estimate the value of $\rho$ at these experiments. Moreover, the measurement of both the event shape and the particle of interest in the same pseudorapidity region induces a bias. The bias can be mitigated by choosing the event shape estimator in the forward rapidity region, which ensures it is in a different pseudorapidity region compared to the particle of interest. Most of the present detectors can measure particle multiplicity in the forward rapidity. Thus, a redefinition of flattenicity is necessary to account for the issues mentioned above, which would consider the charged-particle multiplicity in each cell at the forward rapidity instead of $p_{\rm T}^{\rm cell}$. Flattenicity ($\rho_{\rm ch}$) is determined in the forward-rapidity region, where the ($\eta-\phi$) space is divided into ($8\times 8$) cells and charged particles in each cell ‘$i$’ ($N_{\rm ch}^{\rm{cell},i}$) is estimated. Thus, $\rho_{\rm ch}$ can be estimated using the following equation [106].
@@
Here, $\langle N_{\rm ch}^{\rm{cell}}\rangle$ is the mean number of charged particles in the cells. By construction, $\rho_{\rm ch}$ ranges from 0 to 1, where the lower limit,$\rho_{\rm ch}\rightarrow 0$ indicates isotropic events and $\rho_{\rm ch}\rightarrow 1$ indicates jetty events. To be consistent with other event-shape observables, here, throughout the manuscript, we use $(1-\rho_{\rm ch})$ instead of $\rho_{\rm ch}$, where events with $(1-\rho_{\rm ch})\rightarrow 1$ are likely to be isotropic while $(1-\rho_{\rm ch})\rightarrow 0$ indicates to a jetty event.
Figure 20 shows the correlation of charged-particle flattenicity, ($1-\rho_{\rm ch}$), with $N_{\rm mpi}$. $\rho_{\rm ch}$ is measured using the charged particles having $p_{\rm T}>0$ the V0 acceptance region of ALICE, i.e., $-3.7<\eta<-1.7$ and $5.12.8<\eta<$ 5.1, where ($\eta-\phi$) space is divided into a ($8\times 8$) grid. The correlation of $(1-\rho_{\rm ch})$ with $N_{\rm mpi}$ remains almost flat at 0 till $(1-\rho_{\rm ch})<0.6$, then starts to rise rapidly, where most of the events lie.
It is noteworthy to mention that the charged-particle flattenicity distribution depends on the detector acceptance range from which the charged-particle multiplicity is measured. In this review, we discuss the flattenicity distribution that is obtained by using the multiplicity from the V0M acceptance. In Run 3, a preliminary study has been performed using the FV0A detector to study the flattenicity dependence of strangeness production [107]. However, in the future, it would be interesting to investigate charged-particle flattenicity distribution and the studied observables as a function of charged-particle flattenicity, considering the charged-particle multiplicity from the newly built FT0M and MFT forward detectors of ALICE in Run 3. It is worth mentioning that, to make the studies coherent and detector/experiment independent, one needs to employ an unfolding procedure to retrieve a true flattenicity distribution from the folded distributions with the detector effects.
II.9 Flow vectors ($q_{n}$)
In heavy-ion collisions, the initial configuration of the collision overlap region can affect the final state azimuthal distribution of charged particles. In non-central collisions, the collision overlap region is elliptic in the transverse plane, and thus, the final state azimuthal distributions of particles are elliptic, leading to a larger value of elliptic flow in the mid-central collisions as compared to the most-central collisions. Further, the density fluctuations can lead to observations of higher-order flow coefficients. Interestingly, one can make classifications of events based on the azimuthal distribution of particles in the final state using the reduced flow vectors ($q_{n}$), defined as follows [38, 108]:
@@ II.9 Flow vectors ($q_{n}$)
Here, $M$ is the total number of charged hadrons, and $\phi_{j}$ is the azimuthal angle of the $j$th hadron. $n$ denotes the order of the reduced flow vectors. Here, $q_{2}$ can separate events based on the elliptic distribution of the hadrons in the transverse plane. Similarly, $q_{3}$ isolates events based on the triangular distribution of the hadrons in the transverse plane. For $M\rightarrow\infty$, the reduced flow vectors are proportional to anisotropic flow coefficients of similar order [108]. However, this proportionality becomes weaker as $M$ is reduced [108]. Thus, $q_{n}$ are not good observables for the low multiplicity events [108].
II.10 Aplanarity ($A$)
Aplanarity is another interesting event shape observable that characterizes the shape of the underlying event (UE) based on the momentum of the charged hadrons. Aplanarity is estimated from the linearized sphericity tensor $S^{\rm\mu\nu}$, and defined as  [101, 109, 44];
@@ II.10 Aplanarity ($A$)
  • Aplanarity: $A$ = $\frac{3}{2}$$\lambda_{3}$ measures the amount of transverse momentum in or out of the plane formed by the two leading order eigenvectors, $\lambda_{1}$ and $\lambda_{2}$. Theoretically, $A$ lies between $20\leq A<1/$ 2; however, the typical measured values fall between $0.30\leq A<$ 0.3. The value of $A$ close to zero indicates a relatively planar event, while $A\rightarrow 1/2$ indicates the isotropic events.
  • Sphericity: S = $\frac{3}{2}(\lambda_{2}+\lambda_{3})$ quantifies the total transverse momentum with respect to the sphericity axis defined by the four momenta used in event shape measurement (usually, the first eigenvector). The theoretical values fall between $10\leq S<$ 1. However, the inclusion of the smallest eigenvalue $\lambda_{3}$, the typical maximum value achieved in the experiment is close to S $\simeq 0.8$. S0absent0$\rightarrow$ 0→ 0 indicates a dijet kind of event, while S1absent1$\rightarrow$ 1→ 1 indicates the isotropic events.
@@
  • $C$ = 3($\lambda_{1}\lambda_{2}+\lambda_{1}\lambda_{3}+\lambda_{2}\lambda_{3}$) measures the events with three jets. C0absent0$\rightarrow$ 0→ 0 indicates dijet events.
@@
II.11 Thrust, major and minor, oblatness
@@
$T_{\perp}=\max_{\hat{n}_{\perp}}\frac{|\vec{p}_{\rm Ti}\cdot\hat{n}_{\perp}|}{\sum_{i}|p_{\rm Ti}|}$ (26)
@@
where sum runs over all the hadrons produced in the final state with total momentum $p_{\rm i}$, transverse momentum $p_{\rm Ti}$, and the unit vector $\hat{n}_{\perp}$ is used to maximize the sum. The variable $\tau_{\perp}$ is defined in terms global transverse thrust $T_{\perp}$ in order to keep the common-event shape definition. The allowed values for $\tau_{\perp}$ lies in between the ranges $30<\tau_{\perp}<1/$ 3. This is due to the fact that the allowed values of $T_{\perp}$ fall within the range $30<T_{\perp}<2/$ 3. The smaller the value of $\tau_{\perp}$, denotes for extreme two-jet (i.e cigar-like) events, whereas the higher value denotes spherical events. Here, $T$ is the global event-shape variable thrust. The value of $T$ can vary between 0.5 for spherical events to 1.0 for narrow jet events. Since the thrust variable is linear in momentum, it is more sensitive to low-momentum particles than sphericity.
A plane through the origin and perpendicular to the thrust axis divides the event into two hemispheres $H_{1}$ and $H_{2}$ as shown in Fig. 21. To obtain the thrust major $T_{\rm maj}$, the maximization of Eq. 25 is performed within that plane [111]. Furthermore, the thrust minor $T_{\rm min}$ is given by the argument of Eq. 25 and estimated in the direction perpendicular to both thrust major and thrust axes. It is defined as;
@@
The observable $T_{\rm min}$ is called directly global thrust minor, which measures the out-of-event-plane energy flow.
Using the thrust major $T_{\rm maj}$ and thrust minor $T_{\rm min}$ one can define oblatness ($O$) as [111]
@@
Figure 22: The distribution of thrust $T$, heavy jet mass $\rho$, total and wide-jet broadenings $B_{T}$ and $B_{W}$, $C$ , and $D$ parameters at $\sqrt{s}$ = 91.2 GeV in comparison with various QCD model prediction [113, 114, 115, 116, 117, 118, 119, 78, 120]
Figure 23: The first moment of the event shape variables thrust $T$, heavy jet mass $\rho$, total and wide-jet broadenings $B_{T}$ and $B_{W}$, $C$, and $D$ parameters as a function of center of mass energy $\sqrt{s}$, compared with several QCD model prediction [113, 114, 115, 116, 117, 118, 119, 78, 120]
II.12 Heavy-jet mass ($M_{\rm H}$)
The invariant mass of the particle in each hemisphere, $H_{1}$ and $H_{2}$, is calculated, and the heavy-jet mass $M_{H}$ is defined as the larger of the two [112]. For a given center of mass energy $\sqrt{s}$, one can define the scaled heavy jet mass $\rho_{\rm H}$ as follows;
$\rho=\frac{M_{\rm H}^{2}}{s}$ (30)
II.13 Jet broadening variables $B_{T}$ and $B_{W}$
Jet broadening variables measure of the fraction of energy that is perpendicular to the thrust axis. For each of the two hemispheres $H_{k}$, one can define [121]
$B_{k}=\frac{\sum_{i\epsilon H_{k}}|\vec{p}_{i}\times\hat{n_{\perp}}|}{2\sum_{i}|p_{i}|}.$ (31)
Using the definition of $B_{k}$, two variables are constructed, such that [122]

### v2

@@ 2.8 Charged-particle flattenicity ($\rho_{\rm{ch}}$)
Isotropic events are expected to have a uniform distribution of $p_{\rm T}$ throughout the ($\eta-\phi$) region. To quantify the uniformity of $p_{\rm T}$ distribution event-by-event, the ($\eta-\phi$) region is divided into a $(10\times 10)$ grid. $p_{\rm T}$ in each cell ($p_{\rm T}^{\rm cell}$) can be measured, and corresponding mean ($\langle p_{\rm T}^{\rm cell}\rangle$) and standard deviation ($\sigma_{p_{\rm T}^{\rm cell}}$) can be estimated. Keeping the tracking capabilities of future upgrade for ALICE experiment (ALICE 3) in mind, the charged particles in $|\eta|<4.0$ and $p_{\rm T}>0.15$ GeV/c are considered for the estimation of $\langle p_{\rm T}^{\rm cell}\rangle$ and $\sigma_{p_{\rm T}^{\rm cell}}$. Thus, for an event, flattenicity ($\rho_{\rm ch}$) is defined as the relative standard deviation as follows Ortiz:2022zqr ,
@@
Figure 21: Correlation of flattenicity with number of MPI. The black solid line represents $\langle N_{\rm mpi}\rangle$ as a function of $1-\rho_{\rm ch}$.
Events with both jet signals and underlying event activity are expected to have a larger spread, and correspondingly, $\sigma_{p_{\rm T}^{\rm cell}}$ will be higher. In contrast, events having only soft production of particles will have a narrower distribution of $p_{\rm T}^{\rm cell}$, and therefore corresponding $\sigma_{p_{\rm T}^{\rm cell}}$ will be smaller. It is important to note that the presence of jet signals in the jetty events can also enhance $\langle p_{\rm T}^{\rm cell}\rangle$ in the denominator of Eq. (22); however, the increase in $\langle p_{\rm T}^{\rm cell}\rangle$ will be overshadowed by the corresponding rise in $\sigma_{p_{\rm T}^{\rm cell}}$ in the numerator. Thus, the value of $\rho_{\rm ch}$ is expected to be smaller for isotropic events, while a large value of $\sigma_{p_{\rm T}^{\rm cell}}$ corresponds to a jetty event.
However, the current PID and tracking capabilities at ALICE (LHC) and the STAR experiment (RHIC) are limited to $|\eta|<1.0$. As a result, using the above definition, one can not estimate the value of $\rho_{\rm ch}$ at these experiments. Moreover, the measurement of both the event shape and the particle of interest in the same pseudorapidity region induces a bias. The bias can be mitigated by choosing the event shape estimator in the forward rapidity region, which ensures it is in a different pseudorapidity region compared to the particle of interest. Most of the present detectors can measure particle multiplicity in the forward rapidity. Thus, a redefinition of flattenicity is necessary to account for the issues mentioned above, which would consider the charged-particle multiplicity in each cell at the forward rapidity instead of $p_{\rm T}^{\rm cell}$. Flattenicity ($\rho_{\rm ch}$) is determined in the forward-rapidity region, where the ($\eta-\phi$) space is divided into ($8\times 8$) cells and charged particles in each cell ‘$i$’ ($N_{\rm ch}^{\rm{cell},i}$) is estimated. Thus, $\rho_{\rm ch}$ can be estimated using the following equation Ortiz:2022mfv .
@@
Here, $\langle N_{\rm ch}^{\rm{cell}}\rangle$ is the mean number of charged particles in the cells. By construction, $\rho_{\rm ch}$ ranges from 0 to 1, where the lower limit, $\rho_{\rm ch}\rightarrow 0$ indicates isotropic events and $\rho_{\rm ch}\rightarrow 1$ indicates jetty events. To be consistent with other event-shape observables, here, throughout the manuscript, we use $(1-\rho_{\rm ch})$ instead of $\rho_{\rm ch}$, where events with $(1-\rho_{\rm ch})\rightarrow 1$ are likely to be isotropic while $(1-\rho_{\rm ch})\rightarrow 0$ indicates to a jetty event. This is shown in Fig. 20.
Figure 21 shows the correlation of charged-particle flattenicity, ($1-\rho_{\rm ch}$), with $N_{\rm mpi}$. $\rho_{\rm ch}$ is measured using the charged particles having $p_{\rm T}>0$ the V0 acceptance region of ALICE, i.e., $3.7<\eta<-1.7$ and $2.8<\eta<5.1$, where ($\eta-\phi$) space is divided into a ($8\times 8$) grid. The correlation of $(1-\rho_{\rm ch})$ with $N_{\rm mpi}$ remains almost flat at 0 till $(1-\rho_{\rm ch})<0.6$, then starts to rise rapidly, where most of the events lie.
It is noteworthy to mention that the charged-particle flattenicity distribution depends on the detector acceptance range from which the charged-particle multiplicity is measured. In this review, we discuss the flattenicity distribution that is obtained by using the multiplicity from the V0M acceptance. In Run 3, a preliminary study has been performed using the FV0A detector to study the flattenicity dependence of strangeness production QM2025:flatPoster . However, in the future, it would be interesting to investigate charged-particle flattenicity distribution and the studied observables as a function of charged-particle flattenicity, considering the charged-particle multiplicity from the newly built FT0M and MFT forward detectors of ALICE in Run 3. It is worth mentioning that, to make the studies coherent and detector/experiment independent, one needs to employ an unfolding procedure to retrieve a true flattenicity distribution from the folded distributions with the detector effects.
2.9 Flow vectors ($q_{n}$)
In heavy-ion collisions, the initial configuration of the collision overlap region can affect the final state azimuthal distribution of charged particles. In non-central collisions, the collision overlap region is elliptic in the transverse plane, and thus, the final state azimuthal distributions of particles are elliptic, leading to a larger value of elliptic flow in the mid-central collisions as compared to the most-central collisions. Further, the density fluctuations can lead to observations of higher-order flow coefficients. Interestingly, one can make classifications of events based on the azimuthal distribution of particles in the final state using the reduced flow vectors ($q_{n}$), defined as follows STAR:2002hbo , ATLAS:2015qwl :
@@ 2.9 Flow vectors ($q_{n}$)
Here, $M$ is the total number of charged hadrons, and $\phi_{j}$ is the azimuthal angle of the $j$th hadron. $n$ denotes the order of the reduced flow vectors. Here, $q_{2}$ can separate events based on the elliptic distribution of the hadrons in the transverse plane. Similarly, $q_{3}$ isolates events based on the triangular distribution of the hadrons in the transverse plane. For $M\rightarrow\infty$, the reduced flow vectors are proportional to anisotropic flow coefficients of similar order ATLAS:2015qwl . However, this proportionality becomes weaker as $M$ is reduced ATLAS:2015qwl . Thus, $q_{n}$ are not good observables for the low multiplicity events ATLAS:2015qwl .
2.10 Aplanarity ($A$)
Aplanarity is another interesting event shape observable that characterizes the shape of the underlying event (UE) based on the momentum of the charged hadrons. Aplanarity is estimated from the linearized sphericity tensor $S^{\rm\mu\nu}$, and defined as  Hanson:1975fe , CMS:2018mdd , ALICE:2012cor ;
@@ 2.10 Aplanarity ($A$)
  • 1.
    Aplanarity: $A$ = $\frac{3}{2}$$\lambda_{3}$ measures the amount of transverse momentum in or out of the plane formed by the two leading order eigenvectors, $\lambda_{1}$ and $\lambda_{2}$. Theoretically, $A$ lies between $0\leq A<1/2$; however, the typical measured values fall between $0\leq A<0.3$. The value of $A$ close to zero indicates a relatively planar event, while $A\rightarrow 1/2$ indicates the isotropic events.
  • 2.
[…]
  • 3.
@@
  • 4.
@@
  • 5.
[…]
2.11 Event isotropy
Recently, a novel event shape observable called event isotropy ($\mathcal{I}$) has been designed to identify uniform radiation patterns. Event isotropy is defined as the dimensionless distance of a collider event $\mathcal{E}$ and isotropic radiation pattern $\mathcal{U}$ of the same energy Cesarotti:2020hwb , Cesarotti:2020ngq , ATLAS:2023mny . This distance is evaluated by solving optimal transport problems, using the ‘Energy-Mover’s Distance’ (EMD). It is designed for new physics signals that are far from QCD-like.
$\mathcal{I}\equiv\rm EMD(\mathcal{E},\mathcal{U})$ (27)
The definition of event isotropy depends on the isotropic reference geometries such as spherical, cylindrical, circular (or Rings), Rings with Dipole geometry, and corresponding ground measure Cesarotti:2020hwb , Cesarotti:2020ngq , ATLAS:2023mny . By construction, the event isotropy $\mathcal{I}$ is bounded on $\mathcal{I}$ $\in$ [0, 1]. Perfectly isotropic events take a value of $\mathcal{I}$ = 0, meaning there is zero distance between radiation patterns. Nearly least isotropic (jetty type) events take values approaching $\mathcal{I}$ = 1. To follow the similar historical convention with all other event shape observables, events that the least isotropic (‘dijet-like’) topology is near values of 0 and the most isotropic topology is near values of 1, it is always preferred to use 1- $\mathcal{I}$, instead of $\mathcal{I}$. It is found that event isotropies can differentiate between quasi-isotropic events better than the existing observables $C$ and $D$ parameter Cesarotti:2020hwb , Cesarotti:2020ngq , ATLAS:2023mny .
2.12 Fox-Wolfram moments
In $e^{+}e^{-}$ collisions, a set of unique event shape observables named as Fox-Wolfram moments (FWMs) is introduced to discuss the geometrical shape of the entire jets in an event Fox:1978vu . Later, FWMs are used in collider physics, particularly in high-energy physics, including studies of Higgs and top physics Bernaciak:2012nh and hadron collisions Spiller:2015axa . At LHC energies, FWMs are used to study the medium modifications of jet shapes in relativistic heavy-ion collisions Kong:2024wdk . FWMs, which measure the angular distribution of the energy flow, are characterized by the superposition of spherical harmonics $Y_{l}^{m}$. It is defined as:
$\mathcal{H}_{l}$ $\equiv\left(\frac{4\pi}{2l+1}\right)\sum_{m=-l}^{+l}\left|\sum_{i}Y_{l}^{m}(\Omega_{i})\frac{|\vec{p_{i}}|}{\sqrt{s}}\right|^{2}$
$=\sum_{i,j}\frac{|\vec{p_{i}}||\vec{p_{j}}|}{s}P_{l}(cos\phi_{ij})$ (28)
where the index $i$ and $j$ run over all the hadrons in an event, $\phi_{ij}$ is the angle between particle $i$ and $j$. The energy momentum conservation equation demands $\mathcal{H}_{0}=0$ and $\mathcal{H}_{1}=1$. In practice, $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ are used to describe the geometrical shape of the jets in an event. These moments are sensitive to the number, angular correlation, and energy distribution of jets.
2.13 Thrust, major and minor, oblateness
@@
$T_{\perp}=\max_{\hat{n}_{\perp}}\frac{|\vec{p}_{{\rm T}_{i}}\cdot\hat{n}_{\perp}|}{\sum_{i}|p_{{\rm T}_{i}}|}$ (30)
@@
where the sum runs over all the hadrons produced in the final state with total momentum $p_{\rm i}$, transverse momentum $p_{\rm Ti}$, and the unit vector $\hat{n}_{\perp}$ is used to maximize the sum. The variable $\tau_{\perp}$ is defined in terms global transverse thrust $T_{\perp}$ in order to keep the common-event shape definition. The allowed values for $\tau_{\perp}$ lies in between the ranges $0<\tau_{\perp}<1/3$. This is due to the fact that the allowed values of $T_{\perp}$ fall within the range $0<T_{\perp}<2/3$. The smaller values of $\tau_{\perp}$ denote extreme two-jet (i.e, cigar-like) events, whereas the higher values denote spherical events. Here, $T$ is the global event-shape variable thrust. The value of $T$ can vary between 0.5 for spherical events to 1.0 for narrow jet events. Since the thrust variable is linear in momentum, it is more sensitive to low-momentum particles than sphericity.
A plane through the origin and perpendicular to the thrust axis divides the event into two hemispheres $H_{1}$ and $H_{2}$ as shown in Fig. 22. To obtain the thrust major $T_{\rm maj}$, the maximization of Eq. 29 is performed within that plane Barber:1979yr . Furthermore, the thrust minor $T_{\rm min}$ is given by the argument of Eq. 29 and estimated in the direction perpendicular to both thrust major and thrust axes. It is defined as;
@@
The observable $T_{\rm min}$ is called the directly global thrust minor, which measures the out-of-event-plane energy flow.
Using the thrust major $T_{\rm maj}$ and thrust minor $T_{\rm min}$ one can define oblatness ($O$) as Barber:1979yr
@@
Figure 23: The distribution of thrust $T$, heavy jet mass $\rho$, total and wide-jet broadenings $B_{T}$ and $B_{W}$, $C$ , and $D$ parameters at $\sqrt{s}$ = 91.2 GeV in comparison with various QCD model prediction L3:1992nwf , L3:1992btq , L3:1995eyy , L3:1997bxr , L3:1997dkv , L3:1998ubl , L3:2000shd , L3:2002oql , L3:2004cdh
Figure 24: The first moment of the event shape variables thrust $T$, heavy jet mass $\rho$, total and wide-jet broadenings $B_{T}$ and $B_{W}$, $C$, and $D$ parameters as a function of center of mass energy $\sqrt{s}$, compared with several QCD model prediction L3:1992nwf , L3:1992btq , L3:1995eyy , L3:1997bxr , L3:1997dkv , L3:1998ubl , L3:2000shd , L3:2002oql , L3:2004cdh
2.14 Heavy-jet mass ($M_{\rm H}$)
The invariant mass of the particle in each hemisphere, ${\rm H}_{1}$ and ${\rm H}_{2}$, is calculated, and the heavy-jet mass $M_{\rm H}$ is defined as the larger of the two Clavelli:1981yh . For a given center of mass energy $\sqrt{s}$, one can define the scaled heavy jet mass $\rho_{\rm H}$ as follows;
$\rho_{\rm H}=\frac{M_{\rm H}^{2}}{s}$ (34)
The experiments at LHC incorporate another variant of the jet mass variable which uses the invariant mass ($M_{{\rm H}_{k}}$) of the jet constituents in the hemisphere, ${\rm H}_{k}$ and the scalar sum of the momenta of both sides of the hemisphere, $P$ CMS:2014tkl , CMS:2018svp . Thus, the new definition of the jet mass is given as follows:
$\rho_{{\rm H}_{k}}=\frac{M_{{\rm H}_{k}}^{2}}{P^{2}},$ (35)
where $k=1$ and 2 stand for either of the hemispheres. One can define the total jet mass of an event as follows:
$\rho_{\rm tot}=\rho_{{\rm H}_{1}}+\rho_{{\rm H}_{2}}.$ (36)
A transverse component of the total jet mass can also be calculated in a similar manner, which is denoted as $\rho_{\rm tot}^{\rm T}$.
It is important to note that several studies incorporate jet grooming techniques to selectively remove soft wide-angle radiation within jets that is less likely to have originated from the hard parton that initiated the jet Larkoski:2017jix , Kogler:2018hem . These grooming techniques thereby enable a more meaningful heavy jet mass. Some of the popular grooming techniques include trimming Krohn:2009th , pruning Ellis:2009su , Ellis:2009me , filtering or mass drop tagger (MDT) Butterworth:2008iy , modified MDT Dasgupta:2013ihk , and soft drop Larkoski:2014wba .
2.15 Jet broadening variables $B_{\rm T}$ and $B_{\rm W}$
Jet broadening variables measure the fraction of energy that is perpendicular to the thrust axis. For each of the two hemispheres ${\rm H}_{k}$, one can define Kittel:2005fu
$B_{k}=\frac{\sum_{i\in{\rm H}_{k}}|\vec{p}_{i}\times\hat{n}_{\perp}|}{2\sum_{i}|\vec{p}_{i}|}.$ (37)
Using the definition of $B_{k}$, two variables are constructed, such that Catani:1992jc